# A Hybrid Quantum Neural Network for Split Learning

### v1

@@ Abstract
Quantum Machine Learning (QML) is an emerging field of research with potential applications to distributed collaborative learning, such as Split Learning (SL). SL allows resource-constrained clients to collaboratively train ML models with a server, reduce their computational overhead, and enable data privacy by avoiding raw data sharing. Although QML with SL has been studied, the problem remains open in resource-constrained environments where clients lack quantum computing capabilities. Additionally, data privacy leakage between client and server in SL poses risks of reconstruction attacks on the server side. To address these issues, we propose Hybrid Quantum Split Learning (HQSL), an application of Hybrid QML in SL. HQSL enables classical clients to train models with a hybrid quantum server and curtails reconstruction attacks. In addition, we introduce a novel qubit-efficient data-loading technique for designing a quantum layer in HQSL, minimizing both the number of qubits and circuit depth. Experiments on five datasets demonstrate HQSL’s feasibility and ability to enhance classification performance compared to its classical models. Notably, HQSL achieves mean improvements of over 3% in both accuracy and F1-score for the Fashion-MNIST dataset, and over 1.5% in both metrics for the Speech Commands dataset. We expand these studies to include up to 100 clients, confirming HQSL’s scalability. Moreover, we introduce a noise-based defense mechanism to tackle reconstruction attacks on the server side. Overall, HQSL enables classical clients to collaboratively train their models with a hybrid quantum server, leveraging quantum advantages while improving model performance and security against data privacy leakage-related reconstruction attacks.
@@ 1 Introduction
We present a novel (to the best of our knowledge, the first) application of HQNN to SL, which we call Hybrid Quantum Split Learning (HQSL), and study its security concerning data privacy leakage. In a simple HQSL setup, the HQNN model is divided into two parts: the client-side model, which has only classical neural network layers, and the server-side model, which consists of a quantum layer followed by classical neural network layers. The quantum layer consists of a quantum circuit that converts classical inputs into quantum states, performs quantum computations, and outputs classical data. Considering the limitations of NISQ computers which restrict the number of qubits and circuit size, we employ a low-depth 2-qubit quantum circuit within the quantum layer. This also ensures our simulations have reasonable runtimes. To allow a relatively larger dimension of classical data inputs into our quantum circuit, we propose a qubit-efficient data-loading technique.
To tackle data privacy leakage in HQSL, we integrate a noise layer based on Laplacian distribution into the client-side model. In contrast to existing works on the Laplacian noise layer, we tune the noise parameters considering the rotational properties of the quantum layer’s encoding gates. This paper is also the first, to the best of our knowledge, to study reconstruction attacks in the hybrid quantum-classical domain.
@@
    Feasibility and performance of HQSL with a single client: We propose a novel HQSL architecture by combining SL with HQNN. This allows resource-limited classical clients to collaboratively train HQNN models with a hybrid quantum server featuring a 2-qubit quantum circuit layer followed by classical dense layers. Despite the small size of the quantum circuit, we successfully showcase a modest improvement in the test accuracy and F1-score over the classical counterpart of HQSL by conducting extensive experiments with five standard datasets as a proof of concept. This demonstrates the feasibility and superior performance of HQSL over SL.
@@
    Security aspect of HQSL: We study the security aspect of HQSL against reconstruction attacks on the server side and propose a quantum encoding gate-based Laplacian noise layer defense mechanism as a countermeasure. A comparative study on reconstruction attacks in hybrid and classical settings demonstrates the higher resilience of HQSL compared to SL.
@@
Section 6 Strengthening the security of HQSL against reconstruction attacks, proposed defense mechanism, experiments to test defense mechanism
Section 7 Conclusion
@@ 2 Related Work
Combining Classical and Quantum Resources. Currently available quantum computers exhibit noise, due to numerous factors, e.g., qubit decoherence and gate errors, constraining the complexity of quantum algorithms. Thus, a common practice to leverage the power of quantum computations in the NISQ era is by combining it with classical methods. Hybrid QML models, e.g., Hybrid Quantum Neural Network (HQNN), are becoming a common area of research, where a quantum computer computes only part of the model. For instance, in HQNNs, classical and quantum computing nodes are concatenated, with the classical layers reducing high dimensional data to low dimensions to suit small quantum circuits that make up the quantum layer [26, 2]. This approach is particularly important for current-generation quantum computers, which are constrained by quantum circuit size and depth. Despite the numerous works involving HQNNs, their use and application in a distributed learning domain remain under-explored.
@@
Split Learning (SL) and the Data Privacy Leakage in SL. In the distributed learning domain, SL has addressed computational requirement problems with Federated Learning (FL). In FL, each client has to run the entire ML model, which can particularly be too computationally intensive for resource-constrained clients [34]. By offloading the compute-intensive parts of the model, SL enables more efficient resource allocation on the client devices, benefiting resource-constrained clients.
@@
However, while SL prevents the sharing of clients’ raw data with the server, information leakage still occurs from smashed data [^2] In split learning, smashed data refers to the intermediate output or activations produced by the last layer of the client-side portion of a split neural network.. This is a concerning security issue because the smashed data contain latent information about the raw input data and can be used to stage an input data reconstruction attack [1] at the server side. In the literature, several countermeasures have been proposed against data privacy leakage in SL. Joshi et al. [8] proposed increasing the number of layers in the client-side model. This reduces the mutual information, hence reducing the similarity between the smashed data and raw data. The downside of this technique is that it increases the computational overhead on the client side, which can be problematic for lightweight devices. NoPeek, an information-theoretic measure to reduce information leakage, is another method suggested [36] to reduce data privacy leakage. They do so by augmenting the classification loss function with the distance correlation function to reduce information leakage. However, a privacy-utility tradeoff is a possible consequence of NoPeek. Other methods involve differential privacy [7] and homomorphic encryption, which come at the cost of model performance degradation [29] and too computationally intensive, respectively. In this work, we investigate the problem of data privacy leakage in the hybrid quantum domain and introduce a Laplacian noise-based defense mechanism designed considering the rotational properties of single-qubit rotational gates.
@@ 4.2 Hybrid Quantum Split Learning Model
While several SL configurations have been proposed [35], this work considers the fundamental vanilla SL setup. In this setup, the labels and smashed data are transmitted to the server-side model at the split layer, while the raw data remain on the client side. During backpropagation, the gradients are transmitted from the server side to the client side across the split layer. In HQSL, the quantum layer is introduced as the first layer of the server-side model of SL. The quantum layer consists of a quantum circuit with classical features as input and output. The classical inputs are encoded to quantum states and processed by the quantum circuit. Expectation measurements of the resultant quantum states for each qubit lead to classical features output from the quantum layer (Eq. 3). These are then fed to the next classical layer. The quantum circuit is designed to be of practical importance in the near-term quantum computing era. The general structure of HQSL is shown in Fig. 1. Next, we discuss the construction and selection of the quantum circuit used across this work.
Figure 1: Hybrid Quantum Split Learning (HQSL) model with N clients and 1 hybrid quantum server. The clients have a classical model (assumed to be the same for all clients), and the server model has a quantum layer as its first layer (layer with purple units)
@@ 4.2.1 Construction and Selection of Quantum Circuit
Figure 2: (a) Our qubit-efficient data-loading scheme consists of layers of encoding and parameterized (Param) gates. (b) The quantum circuit that we utilize in this work has RX gates serving as data-loading points (encoding layer). It is designed to make efficient use of qubits while loading data to the quantum circuit
To construct the quantum circuit in HQSL’s quantum layer, we consider the limitations of current-generation quantum computers (see Appendix A for details). Specifically, the circuit width (number of qubits) and depth (number of gates per qubit) are critical considerations that need to be kept as small as possible. The presence or absence of entangling layers is also another important consideration. We trialed 6 different quantum circuit configurations by adjusting the number of qubits, entangling gates, and encoding type. Each configuration was tested by incorporating the circuit as a quantum layer in our HQSL model. The quantum circuit was selected based on the comparison of accuracy and F1-score between HQSL and its classical analogue. We chose the circuit that consistently outperformed SL in all our experiments. We provide the details of these experiments in Appendix C.
The quantum circuit with the optimal performance is shown as Circuit 6 in Fig. 2(b). The circuit consists of 2 qubits with 3-dimensional classical input features $(x_{1},x_{2},x_{3})$ each encoded at 3 different points along each qubit using RX-gates. We drew inspiration from the work done by [25] on the data re-uploading mechanism to design this circuit for this work. The data re-uploading technique involves multiple repetitions of a layer, $L_{i}$, consisting of the generic single-qubit rotation gate, $U\in SU(2)$. The U-gate is used in the encoding layer as $U(\mathbf{X})$ and in the parameterized layer as $U(\mathbf{\Theta})$, where $\mathbf{X}=(x_{1},x_{2},x_{3})$ and $\mathbf{\Theta_{i}}=(\theta_{1}^{i},\theta_{2}^{i},\theta_{3}^{i})$, where $\mathbf{\Theta_{i}}$ represents the parameters for layer $L_{i}$. Thus, each layer can be represented as $L_{i}=U(\mathbf{X})U(\mathbf{\Theta_{i}})$. $L_{i}$, for $i=0,\ldots,N$, is repeated giving rise to the data re-uploading mechanism, and the single-qubit classifier is introduced, given enough repeats, $N$, are performed.
@@
For our work, we adapted the above decomposition of this layer, $L_{i}$, taking into consideration that the width and depth of a circuit should be kept as small as possible, given the limitations due to current quantum devices. Thus, we initially considered a circuit consisting of only a single layer $L$ consisting of 2 qubits, and without any entangling gates. The circuit is represented as Circuit 7 in Fig. 3.
@@
As discussed by [25], entangling gates can improve the classification performance of the circuit. To introduce entanglement within the circuit consisting of a single layer, $L_{i}$, we decomposed the U-gate into RZ–RY–RZ gates and reorganized the circuit such that for each qubit, we uploaded a feature, $x_{k}$ for $k\in\{1,2,3\}$, corresponding to a 3-dimensional input, using rotational gates, followed immediately by a single-qubit parameterized rotational gate of the same type, and a CZ entangling gate between the 2 qubits. This is repeated until all the features have been uploaded. We omit the CZ-gate at the end of the circuit. We show this circuit as Circuit 8 in Fig. 4.
@@
We built both circuits 7 and 8 and tested their performance when utilized in the quantum layer of HQSL. We compared their performance for the classification of the Fashion-MNIST dataset [39] using the accuracy and F1-score metrics – we experimented only on the Fashion-MNIST dataset because it had more room for improvement compared to the other datasets we studied in this paper. All the details for this dataset and these experiments are provided in Section 5. The results, shown in Table 2, demonstrate that the use of entanglement indeed can improve the accuracy and F1-score during the classification of the Fashion-MNIST dataset.
@@
To develop our proposed circuit (Circuit 6), we carried out a final modification to Circuit 8. Specifically, we replaced the embedding gates for loading $\mathbf{X}=(x_{1},x_{2},x_{3})$ into the circuit with RX-gates, as shown in Fig. 2(b). This change in the embedding gates further improved the classification performance, as shown in Table 2.
@@
Table 2: Table comparing 5-fold Mean Accuracy and F1-score of HQSL models consisting of Circuits 7, 8, 6, 9, 10 during classification of Fashion-MNIST dataset. Circuits 9 and 10 are 3-layer versions of Circuit 7 ($L_{i}(j)$) demonstrating the data re-uploading technique with and without entangling CZ-gates respectively ($L_{i}(j)$ represents the $j^{th}$ layer on qubit $i$). Circuit 6, which is a modified version of Circuits 7 and 8, gives the best performance
ID Circuit Mean Accuracy Mean F1-score
7 [Uncaptioned image] 0.812 0.815
8 [Uncaptioned image] 0.823 0.825
6 [Uncaptioned image] 0.839 0.840
9 [Uncaptioned image] 0.824 0.826
10 [Uncaptioned image] 0.821 0.823
@@
The last two rows of Table 2 show the performance of when we adopt the vanilla data re-uploading technique in a 2-qubit circuit with 3 layers, $L$, with (Circuit 9) and without entanglement (Circuit 10), respectively. We used the same entanglement strategy proposed in [25], i.e., using CZ-gates after each layer, except the last layer. Here, we show that our proposed qubit-efficient data-loading circuit (Circuit 6) provides a better accuracy and F1-score than the 3-layer deep circuits (Circuits 9 and 10) that use the vanilla data re-uploading scheme. Our proposed circuit is ‘qubit-efficient’ as, in theory, it allows us to load infinitely large dimensional data to each qubit (this comes at the cost of an infinitely deep quantum circuit and, hence, an unbounded simulation runtime). However, in this work, we restrict the circuit depth by only considering 3-dimensional data being loaded onto a 2-qubit quantum circuit. By imposing this restriction, we keep our simulations within reasonable runtimes, while still providing high performance during classification.
@@
To demonstrate this, we carried out experiments with Circuit 6 by increasing the number of qubits and circuit depth. The resulting accuracy, and simulation time taken for 5-fold training can be seen in Fig 5. In the figure, ‘shallow’ refers to a low-depth circuit with 3 re-uploading points, while ‘deep’ corresponds to a circuit twice as deep as the shallow one, i.e., 6 data-loading points. Hence, a shallow circuit would require 3-dimensional classical features, while a deep circuit would require a dimension of 6.
@@
We highlight here that although as we increase the number of qubits and circuit depth, accuracy increased marginally, the runtime during simulation increases very rapidly (comparing the shallow circuit with 2 qubits and the deep circuit with 8 qubits, an improvement of only around 1.5% causes the time taken for training to increase by more than 500 hours). We also highlight that the deeper (number of layers) and wider (number of qubits) a circuit is, the more difficult it is to implement on currently available quantum computers. For these reasons, it is desirable to keep our quantum circuit shallow and narrow (3 data-loading points and 2 qubits), while still obtaining high accuracy. Hence, our proposed quantum circuit keeps the training time during simulations within reasonable limits, is more likely to be feasible on near-term devices, and provides a better classification performance than a 3-layer deep circuit that uses the data reuploading technique.

### v2

@@ Abstract
Quantum Machine Learning (QML) is an emerging field of research with potential applications to distributed collaborative learning, such as Split Learning (SL). SL allows resource-constrained clients to collaboratively train ML models with a server, reduce their computational overhead, and enable data privacy by avoiding raw data sharing. Although QML with SL has been studied, the problem remains open in resource-constrained environments where clients lack quantum computing capabilities. Additionally, data privacy leakage between client and server in SL poses risks of reconstruction attacks on the server side. To address these issues, we propose Hybrid Quantum Split Learning (HQSL), an application of Hybrid QML in SL. HQSL enables classical clients to train models with a hybrid quantum server and curtails reconstruction attacks. Additionally, we introduce a novel qubit-efficient data-loading technique for designing a quantum layer in HQSL, minimizing both the number of qubits and circuit depth. Evaluations on real hardware demonstrate HQSL’s practicality under realistic quantum noise. Experiments on five datasets demonstrate HQSL’s feasibility and ability to enhance classification performance compared to its classical models. Notably, HQSL achieves mean improvements of over 3% in both accuracy and F1-score for the Fashion-MNIST dataset, and over 1.5% in both metrics for the Speech Commands dataset. We expand these studies to include up to 100 clients, confirming HQSL’s scalability. Moreover, we introduce a noise-based defense mechanism to tackle reconstruction attacks on the server side. Overall, HQSL enables classical clients to train collaboratively with a hybrid quantum server, improving model performance and resistance against reconstruction attacks.
@@ 1 Introduction
We present a novel (to the best of our knowledge, the first) application of HQNN to SL, which we call Hybrid Quantum Split Learning (HQSL), and investigate its resistance against reconstruction attacks due to data privacy leakage. In a simple HQSL setup, the HQNN model is divided into two parts: the client-side model, which has only classical neural network layers, and the server-side model, which consists of a quantum layer followed by classical neural network layers. This design reflects realistic deployment settings, where clients are typically resource-constrained (e.g., edge or IoT devices) and can only support a small number of lightweight classical layers, while compute-intensive operations and all quantum processing are offloaded to the more capable server. The quantum layer consists of a quantum circuit that converts classical inputs into quantum states, performs quantum computations, and outputs classical data. Considering the limitations of NISQ computers, which restrict the number of qubits and circuit size, we employ a low-depth 2-qubit quantum circuit within the quantum layer. This also ensures our simulations have reasonable runtimes. To allow a relatively larger dimension of classical data inputs into our quantum circuit, we propose a qubit-efficient data-loading technique. To tackle data privacy leakage in HQSL, we integrate a noise layer based on Laplacian distribution into the client-side model. In contrast to existing works on the Laplacian noise layer, we tune the noise parameters considering the rotational properties of the quantum layer’s encoding gates. This paper is also the first, to the best of our knowledge, to study reconstruction attacks in the hybrid quantum-classical domain.
@@
    Feasibility and performance of HQSL with a single client: We propose a novel HQSL architecture by combining SL with HQNN. This allows resource-limited classical clients to collaboratively train HQNN models with a hybrid quantum server featuring a 2-qubit quantum circuit layer followed by classical dense layers. Despite the small size of the quantum circuit, we successfully showcase a modest improvement in the test accuracy and F1-score over the classical counterpart of HQSL by conducting extensive experiments with five standard datasets as a proof of concept. This demonstrates the feasibility and superior performance of HQSL over SL. We further validate the practicality of HQSL deployment by evaluating its performance on multiple real quantum devices and noisy simulators.
@@
    Resistance to reconstruction attacks: We study the resistance of HQSL against reconstruction attacks happening at inference time on the server side and propose a quantum encoding gate-based Laplacian noise layer defense mechanism as a countermeasure. A comparative study on reconstruction attacks in hybrid and classical settings demonstrates the higher resilience of HQSL compared to SL.
@@
Section 6 Enhancing the resistance of HQSL against reconstruction attacks, proposed defense mechanism, experiments to test defense mechanism
Section 7 Conclusion and future work
@@ 2 Related Work
Combining Classical and Quantum Resources. Currently available quantum computers exhibit noise, due to numerous factors, e.g., qubit decoherence and gate errors, constraining the complexity of quantum algorithms. Thus, a common practice to leverage the power of quantum computations in the NISQ era is by combining it with classical methods. Hybrid QML models, e.g., Hybrid Quantum Neural Network (HQNN), are becoming a common area of research, where a quantum computer computes only part of the model. For instance, in HQNNs, classical and quantum computing nodes are concatenated, with the classical layers reducing high dimensional data to low dimensions to suit small quantum circuits that make up the quantum layer [30, 2]. This approach is particularly important for current-generation quantum hardware, which is constrained by the number of available qubits, circuit size and depth. Moreover, in the Generative Adversarial Network (GAN) domain, notable works have proposed the use of quantum circuits with classical neural network layers that aid in pre-processing [19] and post-processing [39] giving rise to hybrid quantum-classical GANs, which further highlight the versatility of quantum resources in hybrid architectures. In this work, we propose, for the first time to the best of our knowledge, yet another combination of quantum and classical resources in the form of an HQNN for the split learning domain.
@@
Split Learning (SL) and the Data Privacy Leakage in SL. In the distributed learning domain, SL has addressed computational requirement problems with Federated Learning (FL). In FL, each client has to run the entire ML model, which can be too computationally intensive for resource-constrained clients [38]. By offloading the compute-intensive parts of the model, SL enables more efficient resource allocation on the client devices, benefiting resource-constrained clients.
@@
However, while SL prevents the sharing of clients’ raw data with the server, information leakage still occurs from smashed data [^2] In split learning, smashed data refers to the intermediate output or activations produced by the last layer of the client-side portion of a split neural network.. This is a concerning security issue because the smashed data contains latent information about the raw input data and can be used to stage an input data reconstruction attack [1] at the server side. In the literature, several countermeasures have been proposed against data privacy leakage in SL. Joshi et al. [9] proposed increasing the number of layers in the client-side model. This reduces the mutual information, hence reducing the similarity between the smashed data and raw data. The downside of this technique is that it increases the computational overhead on the client side, which can be problematic for lightweight devices. NoPeek, an information-theoretic measure to reduce information leakage, is another method suggested [41] to reduce data privacy leakage. They do so by augmenting the classification loss function with the distance correlation function to reduce information leakage. However, a privacy-utility tradeoff is a possible consequence of NoPeek. Other methods involve differential privacy [7] and homomorphic encryption, which come at the cost of model performance degradation [33] and too computationally intensive, respectively. In this work, we investigate the problem of data privacy leakage in the hybrid quantum split learning domain and propose a novel noise-based defense mechanism designed considering the rotational properties of single-qubit encoding gates. The methodology, experiments, and results are presented in Section 6.
@@ 4.1 Gate-based Quantum Circuits
The three steps described above (encoding of classical data into quantum states, parameterization via an ansatz, and measurement) form the foundation of Parameterized (or Variational) Quantum Circuits (PQC or VQC). Recent works have explored PQCs in various advanced configurations and structures, such as Block-Ring topologies for enhanced expressivity and entangling capacity [18], and quantum convolutional structures for classification and code recognition tasks [44]. However, for its practicality in the NISQ generation, this work investigates a small-scale quantum circuit featuring a novel data-loading technique, illustrated in Fig. 2 and described in detail in the next section.
@@ 4.2 Hybrid Quantum Split Learning Model
While several SL configurations have been proposed [40], this work considers the fundamental vanilla SL setup. In this setup, the labels and smashed data are transmitted to the server-side model at the split layer, while the raw data remain on the client side. During backpropagation, the gradients are transmitted from the server side to the client side across the split layer (see Fig. 1). In HQSL, the quantum layer is introduced as the first layer of the server-side model of SL. The quantum layer consists of a quantum circuit with classical features as input and output. The classical inputs are encoded to quantum states and processed by the quantum circuit. Expectation measurements of the resultant quantum states for each qubit lead to classical features output from the quantum layer (Eq. 3). These are then fed to the next classical layer on the server-side model. The quantum circuit is designed to be of practical importance in the near-term quantum computing era. A general structure of HQSL is shown in Fig. 1. Next, we discuss the construction and selection of the quantum circuit used in this work.
Figure 1: General structure of a Hybrid Quantum Split Learning (HQSL) model with N clients and 1 hybrid quantum server. The clients have a classical model (assumed to be the same for all clients), and the server model has a quantum layer as its first layer (layer with purple units). The black, green, and orange units can be replaced by any classical neural network component (e.g., convolutional layers) to suit specific classification problems
@@ 4.2.1 Construction and Selection of Quantum Circuit
Figure 2: (a) Our qubit-efficient data-loading scheme consists of alternate layers of encoding and parameterized (Param) gates. (b) The quantum circuit that we utilize in this work has RX gates serving as data-loading points (encoding layer). It is designed to make efficient use of qubits while loading data to the quantum circuit
To construct the quantum circuit in HQSL’s quantum layer, we consider the limitations of current-generation quantum computers (see Appendix A for details). Specifically, the circuit width (number of qubits) and depth (maximum number of gates per qubit) are critical considerations that need to be kept as small as possible. The presence or absence of entangling gates is also another important consideration. We trialed 6 different quantum circuit configurations by adjusting the number of qubits, entangling gates, and encoding type. Each configuration was tested by incorporating the circuit as a quantum layer in our HQSL model. The quantum circuit was selected based on the comparison of accuracy and F1-score between HQSL and its classical analogue. We chose the circuit that consistently outperformed SL in all our experiments. We detail these experiments in Appendix C.
The quantum circuit with the optimal performance is shown as Circuit 6 in Fig. 2(b). The circuit consists of two qubits, each receiving a 3-dimensional classical input feature vector $\mathbf{X}=(x_{1},x_{2},x_{3})$. These features are encoded into each qubit at three distinct points along the circuit using RX-gates. This encoding strategy is inspired by the data re-uploading technique proposed by [29], where classical information is introduced multiple times (using U-gates) throughout the circuit layers to enhance representational capacity. However, our designed circuit differs in two fundamental ways by: (i) encoding input data $\mathbf{X}\in\mathbb{R}^{n}$ at $n$ loading points ($n=3$ for our case) using single-qubit RX-rotations alternating between single-qubit parameterized gates without using U-gates, and (ii) avoiding layer repetitions (re-uploading) to achieve lower circuit depth and fewer parameters while improving performance (see Table 2). We provide more details on these differences and how we designed our circuit next.
The data re-uploading technique involves multiple repetitions of a layer, $L_{i}$, consisting of the generic single-qubit rotation gate, $U\in SU(2)$. The U-gate is used in the encoding layer as $U(\mathbf{X})$ and in the parameterized layer as $U(\mathbf{\Theta})$, where $\mathbf{X}=(x_{1},x_{2},x_{3})$ and $\mathbf{\Theta_{i}}=(\theta_{1}^{i},\theta_{2}^{i},\theta_{3}^{i})$, where $\mathbf{\Theta_{i}}$ represents the parameters for layer $L_{i}$. Thus, each layer can be represented as $L_{i}=U(\mathbf{X})U(\mathbf{\Theta_{i}})$. $L_{i}$, for $i=1,\ldots,N$, is repeated, giving rise to the data re-uploading mechanism, and the single-qubit classifier is introduced, given enough repeats, $N$, are performed.
@@
For our work, we adapted the above decomposition of this layer, $L_{i}$, taking into consideration that the width and depth of a circuit should be kept as small as possible, given the limitations due to current quantum devices. We considered a circuit consisting of only a single layer $L$ consisting of 2 qubits, and without any entangling gates. This circuit has a depth of 6 and consists of a total of 6 trainable parameters ($\theta_{1},\ldots,\theta_{6}$). We represent this circuit as Circuit 7 in Fig. 3.
@@
As discussed by [29], entangling gates can improve the classification performance of the circuit. To introduce entanglement within the circuit consisting of a single layer, $L_{i}$, we decomposed the U-gate into RZ–RY–RZ gates and reorganized the circuit such that for each qubit, we uploaded a feature, $x_{k}$ for $k\in\{1,2,3\}$ using rotational gates, followed immediately by a single-qubit parameterized rotational gate of the same type, and a CZ entangling gate between the 2 qubits. This is repeated until all the features have been uploaded. We omit the CZ gate at the end of the circuit. Hence, this circuit still has a total of 6 trainable parameters, but has now a depth of 8. We show this circuit as Circuit 8 in Fig. 4.
@@
We built both circuits 7 and 8 and tested their performance when utilized in the quantum layer of HQSL. We compared their performance for the classification of the Fashion-MNIST dataset [45] using the accuracy and F1-score metrics–we experimented only on the Fashion-MNIST dataset because it had more room for improvement compared to the other datasets we studied in this paper. All the details for this dataset and these experiments are provided in Section 5. The results, shown in Table 2, demonstrate that the use of entanglement indeed can improve the accuracy and F1-score during the classification of the Fashion-MNIST dataset at the expense of 2 additional CZ gates, increasing the depth by 2, but keeping the number of trainable parameters to 6.
@@
To develop our proposed circuit (Circuit 6), we carried out a final modification to Circuit 8. Specifically, we replaced the embedding gates for loading $\mathbf{X}=(x_{1},x_{2},x_{3})$ into the circuit with RX-gates, as shown in Fig. 2(b). We note that this modification kept the number of trainable parameters and circuit depth of our proposed circuit similar to Circuit 8. This change to RX gates for encoding further improved the classification performance, as shown in Table 2.
@@
Table 2: Table comparing the quantum resource utilization (Number of Trainable Parameters, and Circuit Depth) and 5-fold Mean Accuracy and F1-score of HQSL models consisting of Circuits 7, 8, 6, 9, 10 during classification of Fashion-MNIST dataset. Circuits 9 and 10 are 3-layer versions of Circuit 7 ($L_{i}(j)$), demonstrating the data re-uploading technique with and without entangling CZ gates, respectively ($L_{i}(j)$ represents the $j^{th}$ layer on qubit $i$). Circuit 6, which is a modified version of Circuits 7 and 8, achieves $1-2\%$ accuracy and F1-score improvements with fewer parameters and more than half the circuit depth of standard data re-uploading circuits 9 and 10
ID Circuit Number of Trainable Circuit Mean Mean
Parameters Depth Accuracy F1-score
7 [Uncaptioned image] 6 6 0.812 0.815
8 [Uncaptioned image] 6 8 0.823 0.825
6 [Uncaptioned image] 6 8 0.839 0.840
9 [Uncaptioned image] 18 20 0.824 0.826
10 [Uncaptioned image] 18 18 0.821 0.823
The last two rows of Table 2 show the performance when we adopt the standard data re-uploading technique in a 2-qubit circuit with 3 repeats of layer $L_{i}$, with (Circuit 9) and without entanglement (Circuit 10), respectively. We used the same entanglement strategy proposed in [29], i.e., using CZ-gates after each layer $L_{i}$, except the last layer. In Table 2, we also provide a comparison of quantum resource utilization in terms of the number of trainable parameters, and circuit depth of our proposed quantum circuit compared to the data re-uploading circuits proposed in [29].
@@
Table 2 shows that our proposed shallow (8-deep with 6 trainable parameters) qubit-efficient data-loading circuit (Circuit 6) provides better accuracy and F1-score than the much deeper circuits (Circuits 9 and 10) that use the standard data re-uploading scheme. By loading features exclusively with one-type of single-qubit gates (RX-gates in our case) and omitting repeated layers, Circuit 6 maintains a shallow depth (8) and minimal trainable parameter count (6), compared to the 3-layer data re-uploading circuits (Circuits 9 & 10), which require 18 parameters and depths of 18–20. Our proposed data loading method allows us to achieve better accuracy and F1-score with fewer parameters, improving the trainability and efficiency of our design compared to the standard data re-uploading circuits. Our proposed circuit 6 is ‘qubit-efficient’, as it provides a method to encode a classical input feature vector, $\mathbf{X}$, of arbitrary dimension onto the same qubit. However, this comes at the cost of increasing circuit depth proportionally with input dimensionality. Hence, in this work, we restrict the circuit depth by only considering 3-dimensional data being loaded onto a 2-qubit quantum circuit. By imposing this restriction, we also keep our simulations within reasonable runtimes, while still providing high performance during classification, with a shallow quantum circuit.
@@
To demonstrate this, we carried out experiments with Circuit 6 by increasing the number of qubits and circuit depth. The resulting accuracy, and simulation time taken for 5-fold training can be seen in Fig 5. In the figure, ‘shallow’ refers to a low-depth circuit with 3 data loading points, while ‘deep’ corresponds to a circuit twice as deep as the shallow one, i.e., 6 data-loading points. Hence, a shallow circuit would require 3-dimensional classical features, while a deep circuit would require a data dimension of 6.
@@
We highlight here that although as we increased the number of qubits and circuit depth, accuracy increased marginally, the runtime during simulation increased very rapidly (comparing the shallow circuit with 2 qubits and the deep circuit with 8 qubits, an improvement of only around 1.5% causes the time taken for training to increase by more than 500 hours). We also highlight that the deeper (maximum number of gates) and wider (number of qubits) a circuit is, the more difficult it is to implement on currently available quantum computers. We also face the risk of the barren plateau phenomenon when the number of trainable parameters is excessively high resulting in vanishing gradients [20]. For these reasons, it is desirable to keep our quantum circuit shallow and narrow (3 data-loading points and 2 qubits), while still obtaining high accuracy. In Section 5.1, we effectively provide initial empirical evidence that HQSL with our proposed quantum circuit can be deployed on currently available quantum hardware, providing performance similar to noise-free simulations. Hence, our empirical findings show that our proposed quantum circuit keeps the training time during simulations within reasonable limits, is more likely to be feasible on near-term devices, and provides a better classification performance than a deep circuit with the data re-uploading technique comprising of 3 repeats of layer $L_{i}$.